OpenStudy (anonymous):

What is the gravitational force exerted on the Moon by the Earth? Use Newton's Universal Law of Gravitation. Let's say M1 is the mass of the Earth (5.97 x 10^24 kg) and M2 is the mass of the Moon (7.349 x 10^22 kg). The distance between the Earth and Moon is 384,400,000 m. These values were taken from your book's appendix. Round off your answer to the hundredths placed. NOTE: " x 10^20 " is assumed. So, for example, if your answer is 1.2345 x 10^20, you would simply enter: 1.23

OpenStudy (anonymous):

I have 5.97x10^24 X 7.349x10^22 divide all of that by the (3.844x10^8)^2? and multiply all of that by G, or 6.674x10^-11

OpenStudy (anonymous):

the formula says it is r^2, so I believe I do have to square that 3.844x10^8, right?

OpenStudy (anonymous):

Yea, looks right.

OpenStudy (anonymous):

but it's been so long since me and math have met, that I don't know how to square that? my calculator is saying 1.4776336x10^17 should I trust that?

OpenStudy (anonymous):

1.98 x10^20 is correct? dang that was difficult for me

OpenStudy (anonymous):

Yea, the part with the squaring is good. You can do it this way \[\ (3.844*10^8)^2 = 3.844^2*(10^8)^2 = 14.776 *10^{16} \]

OpenStudy (anonymous):

The final answer is also good.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem